Across at the Midland, one of our members continued the fitment of various bits of woodwork in Gloucester Trailer 56097. These fit behind the luggage rack and will eventually allow the rack to be secured to the framework.
Meanwhile a window which had clouded over and cracked in the blue and grey 108, and this was replaced.
Matador wipers had provided an air schematic for the 141, as the wipers have never worked right since the vehicle was taken into preservation all those years ago. The fault was traced to a mis-labeling of the air pipes, presumably from mainlines days!
Since our last update, the Cravens vehicle has been water blasted, to remove the paint work from the roof and take it back to the metal. This was partially an experiment to see if this method can be used on other vehicles without damaging the body side paintwork, to cut down on the prep time. So far it seems to work quite well. 56456 is now inside the shed ready for the roof to be painted. Work continued over the weekend to install insulation within the vehicle.
51618 was treated to an exam during the week, to return it to be in sync with 56223. The only item of note was the battery terminals which needed a rather good clean.
The broken windows in the Wickham where attended to, though the new piece of glass put up quite a fight, and will need some more attention next time. This gave the opportunity for a rare shot of the power car without the trailer.
On Saturday the 104 was in service on a driver experience train, unfortunately the 47 failed on shed on Saturday morning, the 127 was pressed into service to cover the last two round trips, a rare outing for it solo!
This weekend lead the unusual appearance of the "Cravens" moving from it's normal home behind the shed, out into the open! It is planned to have the roof of the vehicle shot blasted later on in the week, ready for it to be painted.
Following on from the Thomas event last weekend, the three sets received another clean up. With the floors, windows, and seats all needing attention.
A vacuum hose on 51618 was changed in order to improve high side retention. An issue with the gear changer not accepting certain types of spoons was also investigated.
Work continued on the 108, with the pre-cleaners being detached and cleaned up before re-fitment. The rest of the bodywork has also been polished up below the window level, and what a difference it has made.
Another problem encountered is the step plate on 50454 becoming loose. This has been removed and a new one will be machined for fitting later in the month.